OpenStudy (anonymous):

Find the slant asymptote and the vertical asymptotes of the given function. r(x) =(5x^3 + 5x^2)/(x^2 − 4)

OpenStudy (anonymous):

vertical asympote easy. set denominator = 0 and solve to get \[x=2,x=-2\]

OpenStudy (anonymous):

slant asymptote is what you get when you divide.

OpenStudy (anonymous):

what do you mean? Divide what?

OpenStudy (anonymous):

i mean divide \[5x^3+5x^2\] by \[x^2-4\]

OpenStudy (anonymous):

you get \[\frac{5x^3+5x^2}{x^2-4}=5x+5+\frac{20x+20}{x^2-4}\] so your slant asymptote is \[y=5x+5\]